# General Bounds on Functionals of the Lifetime under Life Table Constraints

### 06 Mar 2026 v1

@@ Abstract
In life insurance, life tables are used to estimate the survival distribution of individuals from a given population. However, these tables only provide survival probabilities at integer ages but no information about the distribution of deaths between two consecutive integer values. Many actuarial quantities, such as variable annuities, are functionals of the lifetime and computing them requires full information about mortality rates. One frequent solution is to postulate fractional age assumptions or mortality rate models, but it turns out that the results of the computations strongly depend on these assumptions, which makes it difficult to generalize them. We hence derive upper and lower bounds of functionals of the lifetime with respect to mortality rates, which are compatible with the observed life table at integer ages. We derive two sets of results under distinct assumptions. In the first, we assume that each mortality trajectory is almost surely consistent with all the given one-year survival probabilities from the table. In the second, we consider a relaxed formulation that allows for deviations of the mortality rates while still being consistent in expectation with the given one-year reference survival probabilities. These distinct yet complementary approaches provide a new robust framework for managing mortality risk in life insurance. They characterize the worst- and best-case contract values over all mortality processes that remain compatible with the observed life-table information, thereby enabling insurers to quantify the impact on prices of deviations of the observed mortality rates from their mortality assumptions/models.
@@ 1 Introduction
In this paper, we propose a framework that avoids both sources of model misspecification, namely, the arbitrary choice of fractional age assumption and the structural assumptions of mortality rate models. Instead, we derive upper and lower bounds for actuarial functionals of the lifetime that are compatible with the observed life table at integer ages. Our approach characterizes all possible mortality rates consistent with these discrete data points, and determines the corresponding optimal values of actuarial quantities such as variable annuities. We present two complementary formulations: a strict setting in which each mortality path must match almost surely all the one-year survival probabilities given by the reference life table, and a relaxed (more realistic) setting in which this constraint only holds in expectation, allowing for mortality trajectories to deviate while still being consistent on average with each one-year tabulated survival probabilities. More precisely, for a fixed maturity $T>0$ and a risk-neutral measure $\mathbb{Q}$, we aim at solving stochastic control problems of the form
@@
for some functional $F$ of the lifetime [^1] Strictly speaking, functional of the mortality rates, but we interchangeably use these terms when it is clear from the context. $\mu=(\mu_{t})_{0\leq t\leq T}$, investment fund $A=(A_{t})_{0\leq t\leq T}$, and interest rates $r=(r_{t})_{0\leq t\leq T}$. The strict setting of Section 4 will define the set of admissible controls $\mathcal{A}$ such as to match almost surely the given survival probabilities over each integer year from the life table. As this strict setting effectively leads to deterministic optimal mortality rates, we then consider in Section 5 a relaxed setting where the admissible set $\mathcal{A}$ is such that this constraint only holds in expectation. However, the resulting control problems then tend to be ill-posed, and similar to Avellaneda and Paras (1996) or Li and Szimayer (2011), we need to further restrict this set $\mathcal{A}$ to bounded controls from below and above by time-dependent functions consistent with the life table. We establish asymptotic results showing that, as the set of bounded controls expands, the solutions to the restricted problems converge to the solution of the unrestricted problems.
These two distinct yet complementary approaches let us derive general bounds for functionals of the lifetime that are compatible with observed life tables, enabling insurers to quantify the worst- and best-case contract prices resulting from deviations of the observed mortality rates from their mortality assumptions/models. This framework is therefore not a pricing engine, but instead provides a novel robust and model-free risk management tool to quantify the sensitivity of life-contingent valuations to within-year mortality uncertainty. We highlight that ’model-free’ refers here to the mortality component: no explicit mortality model is ever imposed. A financial model is still required for valuation, see Section 3, but the bounds derived in Section 4 (strict setting) and 5 (relaxed setting) directly generalize to more complex financial dynamics. Numerical experiments in Section 6 further illustrate and support these findings.
@@ 3 The financial market
We then consider the following financial framework for illustrating purposes, as the contract bounds derived in Section 4 and 5 can be extended directly to other and more complex (Markovian) financial dynamics [^3] This is straightforward by adapting the contract price $C_{t}$ in Section 4 and the coefficients of the HJB equations in Section 5.. The interest rate and stock processes, $\left(r_{t}\right)_{0\leq t\leq T}$ and $\left(S_{t}\right)_{0\leq t\leq T}$ are defined on $(\Omega,\mathcal{F},\mathbb{F}:=(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{Q})$ generated by $\mathbf{W}^{\mathbb{Q}}=\left(W_{t}^{1,\mathbb{Q}},W_{t}^{2,\mathbb{Q}}\right)_{0\leq t\leq T}^{\top}$ :
@@ 6 Numerical experiments
Since our method in Section 4 and 5 does not rely on any specific mortality modeling assumptions, the resulting bounds provided below are the most general one can obtain using the available information (provided that the mortality rates over each one-year period are consistent with the given table). These bounds are therefore of high practical interest for insurers, as they naturally provide a range within which the contract price should lie. Any price outside this range is either not consistent with the life tables or with the assumed financial market. Finally, these results allow the insurer to directly quantify its worst-case loss or best-case gain in case of deviation of the observed mortality from the chosen fractional-age/mortality rate assumption. The results are based on the Belgian female life table.
@@ 6.1. Strict bounds
We then study the strict bounds of the GMIB contract in Figure 6.2. For younger policyholders (smaller initial age $x$), the upper-left plot shows that the insurer’s exposure to deviations of mortality within each one-year period from the chosen fractional-age assumption is moderate, at least for relatively short maturities $T$ (while remaining consistent with the tabulated survival probabilities $\hat{p}_{j}$’s). In contrast, for older policyholders (bottom plot) or for longer maturities, this mortality risk becomes much more significant, as deviations within each one-year period from the fractional-age assumption lead to a stronger impact on the GMIB value. We also observe that the choice among the three fractional-age assumptions (2.9)–(2.11) has virtually no impact on the contract price. All these observations are confirmed for the GMDB bounds in Figure 6.3. Bounds for the GMAB contract are not depicted as they are trivial in the strict pathwise setting, see Proposition 4.3.
@@ 7 Conclusion
This paper addresses a structural source of model risk in life insurance: life tables provide survival probabilities at integer ages but leave the within-year timing of deaths unspecified, even though many life-contingent liabilities (e.g. variable annuities) depend on the full continuous-time lifetime distribution. Rather than postulating a fractional-age assumption (UDD/CFM/Balducci) or a parametric (deterministic or stochastic) mortality rate model, we develop a model-free mortality framework that characterizes the set of admissible mortality intensities compatible with the tabulated one-year survival probabilities and then derives sharp upper and lower bounds for actuarial functionals over that set. We present two distinct yet complementary approaches. In the strict (pathwise) matching regime, admissible mortality trajectories must reproduce each observed one-year survival probability almost surely. Because strict matching is often too restrictive empirically and lead to deterministic optimal mortality rates, we then study a relaxed regime where life-table constraints are matched in expectation. To avoid the control problem to be ill-posed, we introduce a regularized admissible set around a baseline mortality intensity consistent with the table, and solve the resulting robust control problem via a dual formulation with Lagrange multipliers and dynamic programming. Numerical experiments confirm intuitive patterns: (i) expectation-based bounds are wider than almost-sure bounds; (ii) adding intermediate-time constraints tightens bounds substantially for GMIB/GMDB; and (iii) the bound asymmetry depends on whether the payoff is survival- or death-contingent, with the effect diminishing at very old ages as survival probabilities shrink. Overall, the proposed bounds provide a practical, model-free mortality risk-management layer on top of any chosen financial model.
@@ Appendix A Proofs
Author contributions Jean-Loup Dupret: Conceptualization, Formal analysis, Investigation, Methodology, Writing. Edouard Motte: Formal analysis, Investigation, Validation, Visualization, Writing.

### 18 Mar 2026 v2

@@ Abstract
In life insurance, life tables are used to estimate the survival distribution of individuals from a given population. However, these tables only provide survival probabilities at integer ages but no information about the distribution of deaths between two consecutive integer values. This incompleteness is particularly relevant for modern insurance products such as variable annuities, whose payoffs depend jointly on lifetime uncertainty and financial market performance. The valuation of such contracts must therefore be carried out in a joint actuarial-financial framework, as their values depend not only on the full information about mortality rates but also on the interaction between mortality risk, asset dynamics, and embedded guarantees. One frequent solution to this incompleteness is to postulate fractional age assumptions or mortality rate models, but it turns out that the results of the computations strongly depend on these assumptions, which makes it difficult to generalize them. We hence derive upper and lower bounds of hybrid functionals of the lifetime with respect to mortality rates, which are compatible with the observed life table at integer ages and the given financial market. We derive two sets of results under distinct assumptions. In the first, we assume that each mortality trajectory is almost surely consistent with all the given one-year survival probabilities from the table. In the second, we consider a relaxed formulation that allows for deviations of the mortality rates while still being consistent in expectation with the given one-year reference survival probabilities. These distinct yet complementary approaches provide a new robust joint actuarial-financial framework for managing mortality risk in life insurance. They characterize the worst- and best-case contract values over all mortality processes that remain compatible with the observed life-table information and the financial market.
@@ 1 Introduction
Therefore, to avoid these two sources of model incompleteness, authors have considered fractional-age interpolation and safe-side reserving — including the sum-at-risk tradition of Lidstone (1905), Norberg (1985), Linnemann (1993) and Kalashnikov and Norberg (2003); the fractional-age comparisons via stochastic ordering of Hürlimann (1990), Willmot (1997), Jones and Mereu (2000), Frostig (2002), Frostig (2003), Barz and Müller (2012); and the the worst-case techniques of Christiansen (2010), Christiansen and Denuit (2010, 2013). However, these approaches remain fundamentally actuarial in both object and method: it seeks to compare admissible lifetime distributions, construct conservative first-order mortality bases, or derive safe-side / worst-case actuarial reserves from one-year life-table information, typically through stochastic ordering, shape restrictions on mortality rates, Thiele-type reserve equations, and the sign of the sum at risk. Barz and Müller (2012) derive tight bounds for functionals of future lifetimes under an increasing force of mortality, while Christiansen and Denuit (2010) construct conservative first-order bases explicitly for safe-side actuarial calculations; Christiansen and Denuit (2013) extend this logic to single- and multiple-decrement settings and mixed survival/death benefits. What these papers do not provide is a joint financial-market framework: there are no traded assets, no self-financing investment strategy, no hedging argument, and no risk-neutral pricing measure. By contrast, the present paper keeps the same life-table incompleteness problem but shifts it into a market-consistent robust valuation setting, where mortality ambiguity is analyzed jointly with financial dynamics and contract pricing rather than only through actuarial reserve calculations. This, in turn, leads to the application of stochastic control techniques over mortality processes, which, to the best of our knowledge, is new to the literature and allows for a much broader admissible mortality class than the shape-restricted intensities of Barz and Müller (2012), Christiansen and Denuit (2010, 2013) and related work.
We therefore propose a framework that avoids both sources of model misspecification, namely, the arbitrary choice of fractional age assumption and the structural assumptions of mortality rate models, while being robust with modern financial valuation. We derive upper and lower bounds for actuarial functionals of the lifetime that are compatible with the observed life table at integer ages. Our approach characterizes all possible mortality rates consistent with these discrete data points, and determines the corresponding optimal values of hybrid actuarial-financial quantities such as variable annuities. We present two complementary formulations: a strict setting in which each mortality path must match almost surely all the one-year survival probabilities given by the reference life table, and a relaxed (more realistic) setting in which this constraint only holds in expectation, allowing for mortality trajectories to deviate while still being consistent on average with each one-year tabulated survival probabilities. More precisely, for a fixed maturity $T>0$ and a risk-neutral measure $\mathbb{Q}$, we aim at solving stochastic control problems of the form
@@
for some functional $F$ of the lifetime [^1] Strictly speaking, functional of the mortality rates, but we interchangeably use these terms when it is clear from the context. $\mu=(\mu_{t})_{0\leq t\leq T}$, investment fund $A=(A_{t})_{0\leq t\leq T}$, and interest rates $r=(r_{t})_{0\leq t\leq T}$. The strict setting of Section 4 will define the set of admissible controls $\mathcal{A}$ such as to match almost surely the given survival probabilities over each integer year from the life table. As this strict setting effectively leads to deterministic optimal mortality rates, we then consider in Section 5 a relaxed setting where the admissible set $\mathcal{A}$ is such that this constraint only holds in expectation. However, the resulting control problems then tend to be ill-posed, and similar to Avellaneda and Paras (1996), Li and Szimayer (2011), we need to further restrict this set $\mathcal{A}$ to bounded controls from below and above by time-dependent functions consistent with the life table. We establish asymptotic results showing that, as the set of bounded controls expands, the solutions to the restricted problems converge to the solution of the unrestricted problems.
These two distinct yet complementary approaches let us derive general bounds for hybrid functionals of the lifetime that are compatible with observed life tables and with modern financial valuation, enabling insurers to quantify the worst- and best-case contract prices resulting from deviations of the observed mortality rates from their mortality assumptions/models. This framework is therefore not a pricing engine, but instead provides a novel robust and model-free risk management tool to quantify the sensitivity of financial life-contingent valuations to within-year mortality uncertainty. We highlight that ’model-free’ refers here to the mortality component: no explicit mortality model is ever imposed. A financial model is still required for valuation, see Section 3, but the bounds derived in Section 4 (strict setting) and 5 (relaxed setting) directly generalize to more complex financial dynamics. Numerical experiments in Section 6 further illustrate and support these findings.
@@ 3 The financial market
We then consider the following financial framework for illustrating purposes, while emphasizing that the contract bounds derived in Section 4 and 5 can be extended directly to other and more complex (Markovian) financial dynamics [^3] This is straightforward by adapting the contract price $C_{t}$ in Section 4 and the coefficients of the HJB equations in Section 5.. The interest rate and stock processes, $\left(r_{t}\right)_{0\leq t\leq T}$ and $\left(S_{t}\right)_{0\leq t\leq T}$ are defined on $(\Omega,\mathcal{F},\mathbb{F}:=(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{Q})$ generated by $\mathbf{W}^{\mathbb{Q}}=\left(W_{t}^{1,\mathbb{Q}},W_{t}^{2,\mathbb{Q}}\right)_{0\leq t\leq T}^{\top}$ :
@@ 6 Numerical experiments
Since our method in Section 4 and 5 does not rely on any specific mortality modeling assumptions, the resulting bounds provided below are the most general one can obtain using the available information (provided that the mortality rates over each one-year period are consistent with the given table). These bounds are therefore of high practical interest for insurers, as they naturally provide a range within which the contract price should lie. Any price outside this range is either not consistent with the life tables or with the specified financial market. Finally, these results allow the insurer to directly quantify its worst-case loss or best-case gain in case of deviation of the observed mortality from the chosen fractional-age/mortality rate assumption. The results are based on the Belgian female life table.
@@ 6.1. Strict bounds
We then study the strict bounds of the GMIB contract in Figure 6.2.
@@
For younger policyholders (smaller initial age $x$), the upper-left plot shows that the insurer’s exposure to deviations of mortality within each one-year period from the chosen fractional-age assumption is moderate, at least for relatively short maturities $T$ (while remaining consistent with the tabulated survival probabilities $\hat{p}_{j}$’s). In contrast, for older policyholders (bottom plot) or for longer maturities, this mortality risk becomes much more significant, as deviations within each one-year period from the fractional-age assumption lead to a stronger impact on the GMIB value. We also observe that the choice among the three fractional-age assumptions (2.9)–(2.11) has virtually no impact on the contract price. All these observations are confirmed for the GMDB bounds in Figure 6.3. Bounds for the GMAB contract are not depicted as they are trivial in the strict pathwise setting, see Proposition 4.3.
@@ 7 Conclusion
This paper addresses a structural source of model risk in life insurance within a joint actuarial-financial framework: life tables provide survival probabilities only at integer ages, leaving the within-year timing of deaths unspecified, even though many modern life-contingent liabilities (variable annuities) depend on the full continuous-time lifetime distribution. Rather than postulating a fractional-age assumption (UDD/CFM/Balducci) or a parametric (deterministic or stochastic) mortality rate model, we develop a model-free mortality framework that characterizes the set of admissible mortality intensities compatible with the tabulated one-year survival probabilities and the given financial market, yielding sharp upper and lower bounds for hybrid actuarial-financial functionals over that set. We present two distinct yet complementary approaches. In the strict (pathwise) matching regime, admissible mortality trajectories must reproduce each observed one-year survival probability almost surely. Because strict matching is often too restrictive empirically and lead to deterministic optimal mortality rates, we then study a relaxed regime where life-table constraints are matched in expectation. To avoid the control problem to be ill-posed, we introduce a regularized admissible set around a baseline mortality intensity consistent with the table, and solve the resulting robust control problem via a dual formulation with Lagrange multipliers and dynamic programming. Numerical experiments confirm intuitive patterns: (i) expectation-based bounds are wider than almost-sure bounds; (ii) adding intermediate-time constraints tightens bounds substantially for GMIB/GMDB; and (iii) the bound asymmetry depends on whether the payoff is survival- or death-contingent, with the effect diminishing at very old ages as survival probabilities shrink. Overall, the proposed bounds provide a practical, model-free mortality risk-management layer on top of any chosen financial model.